OpenStudy (anonymous):

Find the area given by the curves. y=sqrt(x-1), x-y=1

OpenStudy (anonymous):

ok so to find the area between curves you need to determine which curve is on top of teh other right?

OpenStudy (anonymous):

Yeah, I graphed them and the y=sqrt(x-1) is on top. I also set up the intergral so far, but I'm not sure if it is correct. I got \[A=\int\limits_{1}^{2}[(\sqrt{x-1})-(x-1)]dx\]

OpenStudy (anonymous):

yeah it is right good job

OpenStudy (anonymous):

Okay. The part that is confusing me is when you start to simplify it. How would the antiderivate work for (x-1)^1/2?

OpenStudy (anonymous):

kk so (x-1)^1/2 can be thought of as x^1/2

OpenStudy (anonymous):

you know the antiderivative of x^1/2 is 2/3(x^3/2)

OpenStudy (anonymous):

so the int(x-1)^1/2 is 2/3(x-1)^3/2

OpenStudy (anonymous):



OpenStudy (anonymous):

So there is no rule like a chain rule involved in it?

OpenStudy (anonymous):

I got 3/2 as the answer.

OpenStudy (anonymous):

\[\int\limits f^n(x) f'(x) dx=\frac{ f ^{n+1}(x) }{ n+1 }\]

OpenStudy (anonymous):

nah..chain rule is for deriviatives...I used a u-sub with x-1=u

OpenStudy (anonymous):

Oh alright! Thank you!